# On the Structure and Semantics of Identifier Names Containing Closed Syntactic Category Words

### v1

@@ 6.1.2 Context-Specific Differences in Closed-Category Term Usage
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/akimd/bison C 25b3d0e1a3f97a33615099e4b211f3953990c203
@@
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/apache/arrow C++ 4d566e60c5a12f047e1c0f36b9733a56513a7d43
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/jtablesaw/tablesaw Java 5b27ac45e00396fd6dc30572da48ba5181d81ee2
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/pjreddie/darknet C f6afaabcdf85f77e7aff2ec55c020c0e297c77f9
@@
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/haifengl/smile Java d18696f3f3bb198a8ee4a3c7485942508e59080a
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/spring-projects/spring-boot Java c6045c3111c43bd7b0f99e6c2858bfb2999e358f
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/lightbend/config Java d279841c425e70645feacbdc03e41b3b022ad1f9
@@
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/GNOME/gtk C 17bbeb4bc5556a897198613f872543a9a1b4a240
@@
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/ocornut/imgui C++ 75964a9860d497ce3a4b0c482a093e664971466c
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/openssl/openssl C 29864f2b0f1046177e8048a5b17440893d3f9425
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/jedisct1/libsodium C 72fdf1b86c476e81a3352292962423e76cb141d8
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/apache/parquet-java Java c780ffa09be559ca55feb1c0ba996554dd416dea
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/libjpeg-turbo/libjpeg-turbo C 2a0c86278249e7a3c3429caff24c06a50048d772
Digits Distinguisher × Human-Named Convention GUI Builders / Form Designers https://github.com/qt/qtbase C++ 9a06c8ae8c0c13d9a84ee18768e1c68169c31043
Digits Distinguisher × Human-Named Convention Code Generators / Macro Frameworks https://github.com/jhipster/jhipster-bom Java 82e1ab06812348569b80711fc80568575e533c8e
@@
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/godotengine/godot C++ d00f25ecd92f100aedbab4ef4622b95900e9cf01
@@
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/jMonkeyEngine/jmonkeyengine Java dcb22624b426c46805c1468e90adcd0365cd3083
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/PX4/eigen C++ 7cf1c0179eb0f5499dfc1bffbd229783a7865fe1
@@
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/OpenMathLib/OpenBLAS C 70dff3b84fff151f3e58bf54ce54a5e763f29cac
@@ 6.1.3 Closed-Category Term Usage Across System Domains
Table 13 lists all selected systems, the domain they represent, and the Axial Code that motivated their inclusion. To fit the table, we omitted a few details like system size; this can be found in our open data set (Section 11). We analyzed identifiers drawn from five programming contexts—attributes, parameters, functions, declarations, and class names—across two groups of systems: one curated for domain-specific relevance and one composed of general-purpose projects, which were used to construct the CCID (Table 3). The general-purpose group serves as a baseline, as these systems were not selected based on any particular domain hypothesis. Our underlying assumption is that, if closed-category terms are meaningfully correlated with domain-specific concerns, we will observe statistically significant differences in their usage between these two groups.
For each system, we extracted all identifiers and segmented them using Spiral HuckaSpiral . We then filtered out all terms that are neither digits nor included in our predefined lists of closed-category terms (as defined in Section4). After filtering, we compute the normalized frequency of closed-category term usage by dividing the count of qualifying terms by the system’s total lines of code. To assess whether differences in usage were statistically meaningful, we apply a Mann-Whitney U test to compare the distributions between domain-specific and general-purpose systems.
@@
To explore how these different usage profiles affect our results, we conducted a threshold sweep. At each level, a term had to appear in at least a given proportion of systems to be retained. This allowed us to systematically vary our emphasis between ubiquity (terms common across many systems) and specificity (terms concentrated in a smaller, domain-aligned subset). The results, shown in Figure 2, reveal the strongest distributional divergence at thresholds around 0.6 and 0.8. These peaks suggest that both common and moderately specific terms help distinguish domain-specific systems. By contrast, thresholds between 0.1 and 0.3 yielded little significance, likely reflecting linguistic noise from terms with low usage or ambiguous semantic function.
We repeat the analysis at the level of individual closed-category types (prepositions, determiners, conjunctions, digits) to identify which groups drive the observed differences. As shown in Figure 3, prepositions exhibit consistently strong significance across thresholds, particularly between 0.6 and 0.8. Digits and conjunctions show more variable but still notable divergence, while determiners contribute the weakest and least consistent signal. These trends suggest that domain-specific systems rely more heavily on certain linguistic forms, especially prepositions, to express structural or behavioral distinctions central to their design.
To complement the significance testing, we examine Cliff’s Delta as a non-parametric effect size estimate, plotted in Figure 4. This allows us to assess not only whether closed-category usage differs between system types, but also how strongly. The results show that prepositions and digits increasingly favor domain-specific systems at higher thresholds, reaching small to medium effect sizes. Determiners, by contrast, exhibit weak or even negative effect sizes, suggesting a more general-purpose usage profile. Conjunctions remain close to the negligible–small range, with mild domain skew. These patterns reinforce the idea that domain-specific systems do not just differ in which closed-category terms they use, but in how salient those terms are among their most widely reused identifiers.
@@ 6.2 Summary of RQ2
For RQ2, we examine how closed-category terms correlate with multiple forms of context: (1) source-code-local structure, (2) programming language, and (3) broader system domain. Our findings reveal several consistent trends. First, there is no statistically significant difference in the distribution of closed-category terms across the three programming languages under study, though there are some trends that indicate how they may differ in minor (i.e., non-statistically-significant) ways. Second, source code context plays a significant role: Prepositions and conjunctions are used disproportionately in function names; Digits are significantly positively correlated with parameters and class names while significantly negatively correlated with function names; and Determiners are significantly negatively correlated with class names. These patterns align with the communicative roles uncovered in our Selective Codes, such as the use of prepositions to express behavior or data flow, and digits to distinguish instances or versions.
@@
One of the broader aims of RQ2 is to assess whether closed-category terms are meaningful enough to warrant dedicated study. We argue that their statistically significant correlations with specific code contexts support this aim: such terms appear deliberately and consistently in ways that reflect their natural language functions. While our domain-level comparison relies on predefined lists of closed-category terms, without manual verification of each term’s grammatical role, the results nonetheless suggest that these terms may hold particular communicative importance in domain-specific software. This exploratory finding reinforces the potential value of further study. Supporting the appropriate use of closed-category terms through tools, naming conventions, or educational interventions may ultimately benefit program comprehension and internal quality, particularly in domains where such terms help convey behavioral intent.
@@ 7 Related Work
While there are numerous studies on identifier names, this paper represents one of the few to address closed-category terms, and the only paper to do an in-depth study of their usage in open source systems. We discuss relevant related literature below, and how our work can be improved by, or improve upon, their outcomes.
@@ 7.2 Human-subjects studies
Several studies use human subjects to understand the influence and importance of different characteristics of identifiers. Our work is largely complementary to these studies, as it can be used in conjunction with data from these studies to create/support naming techniques. Reem et al. reem:2021 did a survey of 1100 professional developers, shedding some light on developer preferences and practices with respect to the content of identifier names, including things such as the use of abbreviations and preferred identifier length. Feitelson et al. feitelson:2022 studied how the information content of identifiers named affected their memorability, and concluded that short names that contain focused information are likely optimal. Felienne et al. Felienne:2024 find, among other things, that while instructors agree on the importance of naming, there is disagreement between their teaching practices. Even internally, teachers are generally inconsistent in how they teach and practice identifier naming in the classroom. The results of their study highlight the importance of increasing our formal understanding of naming, which can help increase and support consistency of teaching materials and practices in the classroom.
@@ 7.4 Identifier Type and Name Generation
There are many recent approaches to appraising identifier names for variables, functions, and classes. Kashiwabara et al. Kashiwabara:2014 use association rule mining to identify verbs that might be good candidates for use in method names. Abebe Abebe:2013 uses an ontology that models the word relationships within a piece of software. Saeed et al. Saeed:2023 vectorize methods based on metrics and use the K-Nearst Neighbors algorithm with these vectors, and a large data set of methods, to recommend method names. Allamanis et al. Allamanis:2015 introduce a novel language model called the Subtoken Context Model. There has also been work to reverse engineer data types from identifiers malik2019 ; Hellendoorn:2018 . One thing these approaches have in common is the use of frequent tokens and source code context to try and generate high-quality identifier names (or understand their behavior for the purpose of generating types). There is a lot of work in this subfield, but the contrast to our work remains the same for all of them: These approaches aim to predict strong identifier names based on history. Our approach can help, since an understanding of common naming structures can support filtering out names that are inappropriate based on their grammatical structure; teach AI-based approaches how to optimize the identifiers they generate, or at least avoid using bad grammar structure; or help reverse-engineer the semantics of an identifier name based on its grammatical properties. In addition, automated name generation approaches cannot teach us very much about naming practices on their own, or help us formalize our understanding of strong naming structures, and how those can be taught in a classroom. Thus, our work is novel, and complementary to identifier name generation approaches.
@@ 7.5 Software Ontology Creation Using Identifier Names
Many software word ontologies use meta-data about words to understand the relationship between different words. There is a synergistic relationship between the work we present here and software ontologies, since stronger ontologies can help us generate and study grammar patterns effectively, and the CCID can help construct stronger software word ontologies. In particular, studying closed-category terms helps strengthen the metadata used to generate an ontology that seeks to map how words are related to one another, or code behavior.
@@ 7.6 Identifier Structure and Semantics Analysis
Liblit et al. Liblit06cognitiveperspectives discuss naming in several programming languages and makes observations about how natural language influences the use of words in these languages. Schankin et al. Schankin:2018 focus on investigating the impact of more informative identifiers on code comprehension. Their findings show the advantage of descriptive, compound identifiers over short single-word ones. Hofmeister et al Hofmeister:2017 compared comprehension of identifiers containing words against identifiers containing letters and/or abbreviations. Their results show that when identifiers contained only words instead of abbreviations or letters, developer comprehension speed increased by 19% on average. Lawrie et al. Binkley2006 did a study and used three different “levels” of identifiers. The results show that full-word identifiers lead to the best comprehension compared to the other levels studied. Butler’s work  butler2010exploring extends their previous work on Java class identifiers Butler:2009 to show that flawed method identifiers are also associated with low-quality code according to static analysis-based metrics. These papers primarily study the words found in identifiers and how they relate to code behavior or comprehension rather than word metadata (e.g., PoS).
@@
Høst and Østvold study method names as part of a line of work discussed in Høst’s dissertation hostdissertation . This line of work starts by analyzing a corpus of Java method implementations to establish the meanings of verbs in method names based on method behavior, which they measure using a set of attributes which they define HostLexicon . They automatically create a lexicon of verbs that are commonly used by developers and a way to compare verbs in this lexicon by analyzing their program semantics. They build on this work in hostphrasebook by using full method names, which they refer to as phrases, and augment their semantic model by considering a richer set of attributes. The outcome is that they were able to aggregate methods by their phrases and come up with the semantics behind those phrases using their semantic model, therefore modeling the relationship between method names and method behavior. The phrases they discuss are similar to the general grammar patterns studied in our prior work Newman_GP . They extend this use of phrases by presenting an approach to debug method names Host:2009 . In this work, they designed automated naming rules using method signature elements. They use the phrase refinement from their prior paper, which takes a sequence of PoS tags (i.e., phrases) and concretizes them by substituting real words. (e.g., the phrase $<$verb$>$-$<$adjective$>$ might refine to is-empty). They connect these patterns to different method behaviors and use this to determine when a method’s name and implementation do not match. They consider this a naming bug. Finally, in Hst2011CanonicalMN , Høst and Østvold analyzed how ambiguous verbs in method names makes comprehension of Java programs more difficult. They proposed a way to detect when two or more verbs are synonymous and being used to describe the same behavior in a program; hoping to eliminate these redundancies as well as increase naming consistency and correctness. They perform this detection using two metrics which they introduce, called nominal and semantic entropy. Høst and Østvold’s work focuses heavily on method naming patterns; connecting these to the implementation of the method to both understand and critique method naming.

### v4

@@ 6.1.2 Context-Specific Differences in Closed-Category Term Usage
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/akimd/bison C
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/antlr/antlr4 Java
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/PerMalmberg/libcron C++
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/quartz-scheduler/quartz Java
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/apache/commons-collections Java
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/boostorg/container C++
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/rose-compiler/rose C
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/TinyCC/tinycc C
@@
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/apache/arrow C++
@@
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/haifengl/smile Java
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/jtablesaw/tablesaw Java
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/pjreddie/darknet C
@@
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/GNOME/gtk C
@@
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/lightbend/config Java
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/ocornut/imgui C++
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/spring-projects/spring-boot Java
@@
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/godotengine/godot C++
@@
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/jMonkeyEngine/jmonkeyengine Java
@@
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/OpenMathLib/OpenBLAS C
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/PX4/eigen C++
Digits Distinguisher × Human-Named Convention Code Generators / Macro Frameworks https://github.com/jhipster/jhipster-bom Java
Digits Distinguisher × Human-Named Convention GUI Builders / Form Designers https://github.com/qt/qtbase C++
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/apache/parquet-java Java
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/jedisct1/libsodium C
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/libjpeg-turbo/libjpeg-turbo C
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/openssl/openssl C
@@ 6.1.3 Closed-Category Term Usage Across System Domains
Table 13 lists all selected systems, the domain they represent, and the Axial Code that motivated their inclusion. To fit the table, we omitted a few details like system size; this information can be found in our open data set (Section 11.5). We analyzed identifiers drawn from five programming contexts—attributes, parameters, functions, declarations, and class names—across two groups of systems: one curated for domain-specific relevance and one composed of general-purpose projects, which were used to construct the CCID (Table 3). The general-purpose group serves as a baseline, as these systems were not selected based on any particular domain hypothesis. Our underlying assumption is that, if closed-category terms are meaningfully correlated with domain-specific concerns, we will observe statistically significant differences in their usage between these two groups.
For each system, we extracted all identifiers and segmented them using Spiral HuckaSpiral . We then filtered out all terms that are neither numerals nor included in our predefined lists of closed-category terms (as defined in Section 4). After filtering, we compute the normalized frequency of closed-category term usage by dividing the count of qualifying terms by the system’s total lines of code. To assess whether the differences in usage were statistically significant, we applied a Mann-Whitney U test to compare the distributions between domain-specific and general-purpose systems.
@@
To explore how these different usage profiles affect our results, we conducted a threshold sweep. At each level, a term had to appear in at least a given proportion of systems to be retained. This allowed us to systematically vary our emphasis between ubiquity (terms common across many systems) and specificity (terms concentrated in a smaller, domain-aligned subset). The results, shown in Figure 2, reveal the most substantial distributional divergence at thresholds around 0.6 and 0.8. These peaks suggest that both common and moderately specific terms help distinguish domain-specific systems. By contrast, thresholds between 0.1 and 0.3 yielded little significance, likely reflecting linguistic noise from terms with low usage or ambiguous semantic function.
We repeat the analysis at the level of individual closed-category types (prepositions, determiners, conjunctions, numerals) to identify which groups drive the observed differences. As shown in Figure 3, prepositions exhibit consistently strong significance across thresholds, particularly between 0.6 and 0.8. Numerals and conjunctions show more variable but still notable divergence, while determiners contribute the weakest and least consistent signal. These trends suggest that domain-specific systems rely more heavily on certain linguistic forms, especially prepositions, to express structural or behavioral distinctions central to their design.
To complement the significance testing, we examine Cliff’s Delta as a non-parametric effect size estimate, plotted in Figure 4. This allows us to assess not only whether closed-category usage differs between system types, but also how strongly. The results show that prepositions and numerals increasingly favor domain-specific systems at higher thresholds, reaching small to medium effect sizes. Determiners, by contrast, exhibit weak or even negative effect sizes, suggesting a more general-purpose usage profile. Conjunctions remain close to the negligible–small range, with mild domain skew. These patterns reinforce the idea that domain-specific systems do not just differ in which closed-category terms they use, but in how salient those terms are among their most widely reused identifiers.
@@ 6.2 Summary of RQ2
For RQ2, we examine how closed-category terms correlate with multiple forms of context: (1) source-code-local structure, (2) programming language, and (3) broader system domain. Our findings reveal several consistent trends. First, there is no statistically significant difference in the distribution of closed-category terms across the three programming languages under study, though there are some trends that indicate how they may differ in minor (i.e., non-statistically-significant) ways. Second, source code context plays a significant role: Prepositions and conjunctions are used disproportionately in function names; numerals are significantly positively correlated with parameters and class names while significantly negatively correlated with function names; and Determiners are significantly negatively correlated with class names. These patterns align with the communicative roles uncovered in our Selective Codes, such as the use of prepositions to express behavior or data flow, and numerals to distinguish instances or versions.
@@
One of the broader aims of RQ2 is to assess whether closed-category terms are meaningful enough to warrant a dedicated study. We argue that their statistically significant correlations with specific code contexts support the need for further research: such terms appear deliberately and consistently in ways that reflect their natural language functions. While our domain-level comparison relies on predefined lists of closed-category terms, without manual verification of each term’s grammatical role, the results nonetheless suggest that these terms may hold particular communicative importance in domain-specific software. Supporting the appropriate use of closed-category terms through tools, naming conventions, or educational interventions may ultimately benefit program comprehension and internal quality, particularly in domains where such terms help convey behavioral intent.
@@ 7 Related Work
While numerous studies have been conducted on identifier names, this paper represents one of the few to address closed-category terms, and the only paper to conduct an in-depth analysis of their usage in open-source systems. We discuss relevant related literature below, and how our work can be improved by, or improve upon, their outcomes.
7.1 Grounded Theory in Software Engineering
Our methodology is inspired by Straussian grounded theory CorbinStrauss:1990 , which emphasizes iterative coding, constant comparison, and the development of conceptual categories grounded in data. While we drew inspiration from classic Straussian grounded theory, our approach adapts it to the structure of source code, treating identifiers as theory-generating artifacts rather than unstructured text or human narratives. We employed several key components of the method, including open coding, memoing, axial coding, selective coding, and constant comparison, to analyze identifier names in the context of surrounding code and comments.
Our analysis reached theoretical saturation in that we developed axial codes iteratively until they accounted for all observed behaviors, refining them as new cases emerged. We also engaged in a form of core category development through cross-category synthesis, identifying broader trends that link behavioral interpretations across different grammatical constructs such as determiners, prepositions, digits, and conjunctions.
As grounded theory was originally developed in sociology, its application to the semi-structured nature of software artifacts, including lexical constructs like variable names, presents unique challenges. Some of these challenges have been explored in prior work on grounded theory in software engineering Adolph2011 ; Stol2016 ; hoda2022 , which informed our adapted approach. Our methodology is an example of how grounded-theory-based techniques can support the analysis of naming practices, particularly when grammar patterns and part-of-speech information are involved.
@@ 7.3 Human-subjects studies
Several studies use human subjects to understand the influence and importance of different characteristics of identifiers. Our work is largely complementary to these studies, as it can be used in conjunction with data from these studies to create/support naming techniques. Reem et al. reem:2021 conducted a survey of 1100 professional developers, shedding light on developer preferences and practices regarding the content of identifier names, including the use of abbreviations and preferred identifier length. Feitelson et al. feitelson:2022 studied how the information content of identifiers named affected their memorability, and concluded that short names that contain focused information are likely optimal. Felienne et al. Felienne:2024 find, among other things, that while instructors agree on the importance of naming, there is disagreement between their teaching practices. Even internally, teachers are generally inconsistent in how they teach and practice identifier naming in the classroom. The results of their study highlight the importance of increasing our formal understanding of naming, which can help grow and support the consistency of teaching materials and practices in the classroom.
@@ 7.5 Identifier Type and Name Generation
There are many recent approaches to appraising identifier names for variables, functions, and classes. Kashiwabara et al. Kashiwabara:2014 use association rule mining to identify verbs that might be good candidates for use in method names. Abebe Abebe:2013 uses an ontology that models the word relationships within a piece of software. Saeed et al. Saeed:2023 vectorize methods based on metrics and use the K-Nearst Neighbors algorithm with these vectors, and a large data set of methods, to recommend method names. Allamanis et al. Allamanis:2015 introduce a novel language model called the Subtoken Context Model. There has also been work to reverse engineer data types from identifiers malik2019 ; Hellendoorn:2018 . One thing these approaches have in common is the use of frequent tokens and source code context to try and generate high-quality identifier names (or understand their behavior for the purpose of generating types). There is a lot of work in this subfield, but the contrast to our work remains the same for all of them: These approaches aim to predict strong identifier names based on history. Our approach can help, since an understanding of common naming structures can support filtering out names that are inappropriate based on their grammatical structure; teach AI-based approaches how to optimize the identifiers they generate, or at least avoid using bad grammar structure; or help reverse-engineer the semantics of an identifier name based on its grammatical properties. In addition, automated name generation approaches cannot teach us much about naming practices on their own, nor can they help us formalize our understanding of strong naming structures and how those can be taught in a classroom. Thus, our work is novel, and complementary to identifier name generation approaches.
@@ 7.6 Software Ontology Creation Using Identifier Names
Many software word ontologies use meta-data about words to understand the relationship between different words. There is a synergistic relationship between the work presented here and software ontologies, as stronger ontologies can facilitate the effective generation and study of grammar patterns, and the CCID can aid in constructing stronger software word ontologies. In particular, studying closed-category terms helps strengthen the metadata used to generate an ontology that seeks to map how words are related to one another, or code behavior.
@@ 7.7 Identifier Structure and Semantics Analysis
Liblit et al. Liblit06cognitiveperspectives discuss naming in several programming languages and make observations about how natural language influences the use of words in these languages. Schankin et al. Schankin:2018 focus on investigating the impact of more informative identifiers on code comprehension. Their findings show the advantage of descriptive, compound identifiers over short single-word ones. Hofmeister et al Hofmeister:2017 compared comprehension of identifiers containing words against identifiers containing letters and/or abbreviations. Their results show that when identifiers contained only words instead of abbreviations or letters, developer comprehension speed increased by 19% on average. Lawrie et al. Binkley2006 did a study and used three different “levels” of identifiers. The results show that full-word identifiers lead to the best comprehension compared to the other levels studied. Butler’s work  butler2010exploring extends their previous work on Java class identifiers Butler:2009 to show that flawed method identifiers are also associated with low-quality code according to static analysis-based metrics. These papers primarily study the words found in identifiers and how they relate to code behavior or comprehension rather than word metadata (e.g., PoS).
@@
Høst and Østvold study method names as part of a line of work discussed in Høst’s dissertation hostdissertation . This line of work starts by analyzing a corpus of Java method implementations to establish the meanings of verbs in method names based on method behavior, which they measure using a set of attributes which they define HostLexicon . They automatically create a lexicon of verbs that are commonly used by developers and a way to compare verbs in this lexicon by analyzing their program semantics. They build on this work in hostphrasebook by using full method names, which they refer to as phrases, and augment their semantic model by considering a richer set of attributes. The outcome is that they were able to aggregate methods by their phrases and come up with the semantics behind those phrases using their semantic model, therefore modeling the relationship between method names and method behavior. The phrases they discuss are similar to the general grammar patterns studied in our prior work Newman_GP . They extend this use of phrases by presenting an approach to debug method names Host:2009 . In this work, they designed automated naming rules using method signature elements. They use the phrase refinement from their prior paper, which takes a sequence of PoS tags (i.e., phrases) and concretizes them by substituting real words. (e.g., the phrase $<$verb$>$-$<$adjective$>$ might refine to is-empty). They connect these patterns to different method behaviors and use this to determine when a method’s name and implementation do not match. They consider this a naming bug. Finally, in Hst2011CanonicalMN , Høst and Østvold analyzed how ambiguous verbs in method names makes comprehension of Java programs more difficult. They proposed a method to detect when two or more verbs are synonymous and used to describe the same behavior in a program, aiming to eliminate these redundancies while increasing naming consistency and correctness. They perform this detection using two metrics which they introduce, called nominal and semantic entropy. Høst and Østvold’s work focuses heavily on method naming patterns; connecting these to the implementation of the method to both understand and critique method naming.